The heat produce when a certain mass of brass is heated from 25 ᵒC to 65 ᵒC is 465 Calories. The specific heat of is 0.090. Calculate the mass of brass to produce this requirement.



1
Expert's answer
2021-05-21T10:42:27-0400

"465\ (cal)=1946.86\ (J)"


"0.09\ (cal/(kg\cdot K))=377\ (J/(kg\cdot K))"


"Q=cm\Delta t\to m=Q/(c\Delta t)=1946.86/(377\cdot(65-25))=0.1291\ (kg)" . Answer